# Performance and Generalizability Impacts of Incorporating Geolocation into Deep Learning for Dynamic PM2.5 Estimation

### v1

@@ 1 Introduction
The same principles of spatial dependence, connectivity, and location underlie deep learning (DL) architectures, which have revolutionized many domains in science and now in society under the umbrella brand of artificial intelligence (AI). In fact, the most successful and revolutionary architectures of deep learning are those that successfully leverage spatial context (i.e., spatial dependence) [4, 5], or temporal context [6], or both [7]. These deep learning architectures have found significant performance advantages over point-to-point prediction tasks. In point-to-point prediction, observations of $x_{i}$ at location $i$ are used to predict $y_{i}$ at the same location. Instead, modern deep learning models leverage spatial and spatiotemporal dependencies. To formalize spatial and temporal context in this manuscript, consider the following:
@@
Spatial Context: Given a set of spatial locations $S=\{s_{1},s_{2},\dots,s_{n}\}$, where $s_{i}$ represents the coordinates of location $i$, the observation $x_{i}$ at location $s_{i}$ is influenced by its spatial neighborhood $N_{s}(s_{i})=\{s_{j}:d(s_{i},s_{j})\leq r,j\neq i\}$, where $d(\cdot,\cdot)$ can be a distance metric and $r$ is a threshold radius (or another definition of connectivity or neighborhood). A spatial prediction model can be formulated as:
@@
where $f(\cdot)$ is the prediction function learned by the model. Various flavors of Convolutional Neural Networks, CNNs [8], which belong to this group where the data lie on a regular grid and the spatial neighborhood is the set of pixels in a fixed convolution window, have been widely adopted in geospatial applications and remote sensing for a variety of tasks, for land [9] or ocean remote sensing [10].
Temporal Context: For temporal sequences, consider a discrete time index $T=\{t_{1},t_{2},\dots,t_{T}\}$. The observation $x_{i}^{t}$ at location $i$ and time $t$ is influenced by its past observations $\{x_{i}^{t-k},\dots,x_{i}^{t-1}\}$ for some lag $k$. A temporal prediction model is represented as:
@@
Spatiotemporal Context: Combining both spatial and temporal dimensions, the prediction at location $i$ and time $t$ depends on the spatiotemporal neighborhood:
@@
In this family of models, ConvLSTMs [14] capture spatiotemporal context and have gained adoption in environmental monitoring [15]; Transformers [4] are being employed for remote sensing and geospatial applications [16], and spatiotemporal graph neural networks have shown promise in spatial epidemiology [17], or even in remote sensing for building change detection [18].
@@
The taxonomy above provides a non-exclusive overview of deep learning approaches incorporated in geospatial data science; for instance, sequence-to-sequence methods are not discussed here, but the reader can refer to [19]. The intention is to highlight that the revolutionary impact of deep learning methods is in large due to their ability to successfully capture spatiotemporal context, and how there is potential for further enhancement if this context is made richer, primarily by capturing (geo)location information in addition to spatial information.
Spatial versus Geographic: While advances in deep learning excel at leveraging spatial information, optimal ways of leveraging geographic location information remain underexplored. Note that several methods are proposed to transform geographic (or projected) coordinates into features for deep learning [20], however, regardless of (and beyond) the transformations used, the general higher-level approach in incorporating location into deep learning, which is still an active area of research, is our focus here.
It is crucial that we first differentiate between spatial and geographic. In the context of a CNN, for example, the network learns to identify patterns within the image coordinates system, for instance around point $(x,y)$ in image coordinates, without developing an understanding of geographic location $s_{i}$ corresponding to latitude $\varphi$ and longitude $\lambda$ mapped to $(x,y)$. As mentioned above, [20] surveys a list of methods to input geolocation into deep learning. However, how to best incorporate location (i.e., latitude $\varphi$ and longitude $\lambda$) to maximize generalizability remains an open question, with little research quantifying the generalizability impacts of doing so using different methods on temporally-dynamic estimation scenarios. This is partly due to the gap between studies applying deep learning in geospatial sciences and environmental monitoring, where predictive performance is measured on training and testing sets with minimal interrogation of the black box nature of deep learning, and the scarcity of fundamental studies quantifying the behavior of (geo)location information within deep learning frameworks. Investigating the role of location in deep learning-based estimation is even more important given the emergence of ambitious location encoders [21, 22, 23], which attempt to summarize all relevant information at location $s$ into an embedding vector $e_{s}$, suitable for ingestion by deep learning algorithms, for improved predictive performance, and more importantly, geographic generalizability.
State of the Art: Prior research has made significant progress in identifying strategies to incorporate spatial dependence [12] or geographic information into deep learning, including direct use of raw coordinates, hand-engineered spatial features, and learned location embeddings. Mai et al. [20] provide a comprehensive survey of coordinate transformation methods—such as sinusoidal encodings, spatial tile embeddings, and kernel-based representations. Separately, contrastive learning frameworks such as CSP [21], SatCLIP [22], and GeoCLIP [23] have adapted CLIP-style architectures [24] to align image and location representations in a shared latent space, enabling pretrained models to distill contextual geographic knowledge. However, evaluations of these location encoders have been limited to static tasks such as species distribution or primitive tasks such as predicting long-term climate averages, where temporal dynamics and real-time variability are less critical, and real-world application is limited. On the other hand, studies such as [25, 26, 17] have proposed robust spatial validation—such as checkerboard partitioning or distance-aware cross-validation—to more accurately assess geographic generalizability. Yet, the interplay between location encoding strategies and spatial evaluation methodology remains underexplored, particularly for dynamic, high-resolution prediction tasks.
Contribution: In this paper, we take a step towards closing this gap by quantifying and analyzing the impact of incorporating geolocation information of locations $s_{i}$ with latitude $\varphi$ and longitude $\lambda$ in deep learning within a real-world temporally-dynamic application context of estimating daily surface-level PM2.5 from remote sensing imagery. We first offer an in-depth discussion of different approaches to incorporating geolocation into deep learning, including simple featurization to leveraging the state-of-the-art location encoders. We then characterize the domain of surface-level air pollution estimation, including a brief overview and justification for selecting this application domain as an appropriate test-bed for this study. We then present a series of experiments and results from two complementary aspects of predictive (i.e., estimation) performance and geographic generalizability using Within-Region (WR) and Out-of-Region (OoR) evaluation scenarios. We present an ablation study to compare alternatives, and complement the experimental results with qualitative analysis of wall-to-wall products in the contiguous United States. We characterize the findings and suggest directions for future research.
While the primary contribution of this paper is centered on methods of incorporating geolocation into deep learning, our work also contributes to the literature on surface-level air pollution estimation. Unlike existing PM2.5 estimation models that either rely on point-to-point regression or treat spatial context through static features like land use or proximity to emission sources, our current study advances the field by explicitly interrogating the role of geolocation in deep learning for dynamic, high-resolution air pollution estimation. Our prior work introduced a state-of-the-art Bi-LSTM with attention architecture that demonstrated superior performance, particularly during high-pollution events such as wildfires, by incorporating temporal inputs, including aerosol and meteorological data, and by integrating wildfire smoke density as a predictive covariate [27, 28, 29]. This paper extends that study and evaluates and compares multiple geolocation integration strategies within the modeling framework. By evaluating both within-region (WR) and out-of-region (OoR) performance using spatial partitioning schemes, we move beyond conventional random validation schemes common in air pollution studies. This explicit attention to geographic generalizability and the evaluation of location encoder representations in a dynamic prediction task distinguishes our work from prior efforts, including Di et al. [30], Wei et al. [31], and even our own earlier study [27], which did not examine location feature representations or spatial generalization as central research questions.
2 Geolocation and Deep Learning
As discussed in section 1, deep learning architectures, which were primarily developed on non-geographic data such as natural imagery and text data, excel at leveraging spatial and temporal (sequence) context, but not geographic context. In this section, we broadly describe different methods for incorporating geolocation in deep learning for geospatial applications. In our discussions, we focus on a general case of estimating a target variable $y$ using a set of observations $x_{1}$ through $x_{m}$, with each $x_{j}$ describing a predictive variable observed at location $s_{i}$. For simplicity and staying focused on our core issue, we limit our notations to $s_{i}$, and do not incorporate spatial context of locations $s_{1}$ through $s_{n}$ in $S_{n}$ (spatial neighborhood) or temporal context (in location $s_{i}$) in our notation, given that the strength to incorporate such spatial or temporal context already exists in models such as CNNs, LSTMs, Graph Neural Networks (GNNs), and Transformers. Rather, our focus is on how geolocation, i.e., latitude $\varphi\in[-\pi/2,\pi/2]$ and longitude $\lambda\in[-\pi,\pi]$ of location $s_{i}$ (and by extension, $s_{1}$ through $s_{n}$ in $S_{n}$) can be used in conjunction with observations $x_{1}$ through $x_{m}$ for improved predictive performance as well as geographic generalizability. The same approaches described below for incorporating geolocation can be used with temporal or spatiotemporal models, and in fact, our experiment domain in this paper leverages a spatiotemporal model.
2.1 Approach 1: No Geolocation Feature
@@
2.2 Approach 2: Geolocation as Geographic Coordinates Features
@@
Please note that the focus of this paper is not to discuss a comprehensive list of ways to mathematically transform (i.e. wrap) geolocation as ingestible feature, which has been surveyed in [20]. Rather, we use this representative sinusoidal transformation approach to then quantify impact on performance and more importantly, geographic generalizability, as compared to pre-learned location encoders. Regardless of the transformation method, direct incorporation of geographic location as features has the same impact on generalizability in the contiguous US (our study area), as the latitude and longitude ranges (approximately $24^{\circ}$$49^{\circ}$ and $67^{\circ}$$125^{\circ}$) do not introduce significant discontinuities; whether using raw coordinates, sinusoidal wrapping, or an alternative such as Fourier feature expansion, the generalization outcome remains largely unchanged.
The use of latitude and longitude as model inputs is not novel on its own. Established climate models, such as General Circulation Models (GCMs), simulate the Earth’s climate system by discretizing the globe into latitude-longitude grid cells [32]. In GCMs, each grid cell represents a specific geographic region, and climate variables (e.g., temperature, precipitation) are simulated based on the cell’s location, considering factors such as solar radiation, altitude, proximity to oceans, along with physical and chemical processes. In the same vein, incorporating latitude $\varphi$ and longitude $\lambda$ in deep learning models anchors predictions to specific locations, potentially capturing spatially-varying relationships, just as GCMs account for geographic variation in climate processes.
Including $\varphi$ and $\lambda$ as input features has important practical implications in the context of parameter-heavy and highly non-linear deep learning: the model will parameterize a complex, non-linear mapping between location and the target variable $y$. This is a risky double-edged sword: on the one hand, it can potentially improve performance if the testing region is within the training region. For instance, the winning team of the AutoICE challenge, which was a benchmark-based competition for improving deep-learning-based mapping of sea ice parameters, reports to have used geolocation features to further improve their model, potentially giving them an advantage in the leader board [33]. This advantage was primarily driven by the fact that the testing region was within the training region (East coast of Greenland and the Canadian Archipelago), whose selection by the benchmark designers was originally motivated by ensuring the same sea ice class distribution, but there was no rule to not use geolocation features for training the models [34].
@@
2.3 Approach 3: Geolocation as Pre-learned Location Encodings
@@
A general formulation of the positional encoder is:
@@
3 Experiments
3.1 Domain Characterization: Estimating Surface-level PM2.5
Estimating surface-level air pollution, and in particular, PM2.5 is important to quantify for public health studies. Prolonged exposure to high concentrations of PM2.5 has been linked to respiratory and cardiovascular diseases, and premature mortality [35]. In the United States, the Environmental Protection Agency (EPA) has deployed a series of carefully-calibrated Air Quality System (AQS) monitors to measure the spatially and temporally variable concentrations of PM2.5. However, public health studies require estimation of surface-level pollution at the addresses of study cohort members, while the AQS stations may be far away and spatially-sparse. On the other hand, satellite observations can be used to estimate particulates suspended in columns of air, and therefore, the predictive task is to use satellite observations and ancillary data to derive surface-level (rather than total column) concentrations of PM2.5. Put differently, the machine learning target is the surface-level PM2.5 levels measured by AQS sensors.
@@ 3 Experiments
3.2 Model Architecture
@@
Our base model for PM2.5 concentration prediction (Fig.1(a)) follows a Bidirectional Long Short-Term Memory (Bi-LSTM) [6, 36, 37] network with Luong Attention [38], as detailed in [27]. In brief, a 21-day window of time-series multi-variate input is fed into a Bi-LSTM to capture forward and backward dependencies in PM2.5-related features, which include satellite-derived Aerosol Optical Depth (AOD) [39], meteorological variables (e.g., temperature, precipitation, wind direction/speed) [40, 41], wildfire smoke density [29, 28], elevation [42], Normalized Difference Vegetation Index (NDVI) [43], K-Nearest Neighbors Inverse Distance Weighted (IDW) PM2.5 measurements to incorporate ground-based spatial context, as well as temporal encodings (sine and cosine of (Day of the Year, Month of the Year), and Year).
@@
These features are selected to capture the primary drivers of surface-level PM2.5 concentrations. AOD is satellite-observed columnar aerosol loadings, which, though not a direct measure of surface-level concentrations, is predictive of surface-level pollution under certain atmospheric conditions. Meteorological variables influence pollutant dispersion, chemical transformation, and accumulation. Wildfire smoke density directly relates to episodic spikes in PM2.5 primarily due to biomass combustion. NDVI provides information on vegetative cover, which can modulate emissions and pollutant deposition. Elevation influences local meteorology and pollutant trapping, while the KNN-IDW interpolated PM2.5 layer reflects observed surface-level spatial trends from monitoring stations. Temporal encodings help capture seasonal patterns in emissions.
@@
A masking layer is used to handle any missing data in the inputs [44]. Layer normalization and dropout [45] are applied after each Bi-LSTM layer to stabilize training and mitigate overfitting. The Luong attention mechanism [38] then learns to focus on the most informative timesteps within the 21-day window by computing a context vector as a weighted average of the hidden states. The representation is then passed through a fully connected layer to make the final prediction for PM2.5. We refer the reader to [27] for detailed information on data processing, implementation, long-term evaluation of the model and comparison to multiple other baselines and benchmarks.
@@
Figure 1: Model architecture for estimating surface-level PM2.5 concentrations using a) no geolocation info, or geolocation info as geographic coordinates features, and b) fused geolocation information extracted from the pretrained location encoder.
We evaluate four variations of this model—differing only in the way geolocation is incorporated:
@@
    No geolocation input: The model does not receive geolocation $(\varphi,\lambda)$ information, similar to Approach 1 described in section 2, forcing the model to purely rely on time-series observations for estimating the target.
@@
    Direct geolocation input: We append the raw latitude and longitude values $(\varphi,\lambda)$ to the time-series feature vector at each timestep, similar to Approach 2 described in section 2, allowing the model to rely on both the time series observations as well as geolocation for estimating the target.
@@
    Sinusoidal transformation of geographic coordinates: Also similar to Approach 2 in section 2, but with a transformation applied first: instead of passing latitude and longitude values directly in angular units, we apply sinusoidal transformations to preserve their cyclical properties:
@@
    Fusing location encoder embeddings: We integrate pre-learned location embeddings from the pretrained location encoder GeoCLIP [23]. GeoCLIP is a location encoder trained using a CLIP-style contrastive learning framework to align geotagged Flickr images with geographic coordinates (section 2.3 above covers details of this approach). Its image encoder is a Vision Transformer (ViT-B/16), while the positional encoder is a Fourier feature mapping of latitude and longitude, capturing multi-scale spatial patterns. During pretraining by the original developers [23], GeoCLIP was trained on over 1.2 million image-location pairs globally sampled from the YFCC100M dataset, with relatively denser coverage in urbanized and photo-rich regions, which is important for PM2.5 estimation in areas more prone to generate emissions. We will contextualize GeoCLIP against other available location encoders in section 4.3.

### v2

# Performance and Generalizability Impacts of Incorporating Location Encoders into Deep Learning for Dynamic PM2.5 Estimation

@@ 1 Introduction
The same principles of spatial dependence, connectivity, and location underlie deep learning (DL) architectures, which have revolutionized many domains in science and now in society under the umbrella brand of artificial intelligence (AI). In fact, the most successful and revolutionary architectures of deep learning are those that successfully leverage spatial context (i.e., spatial dependence) [4, 5], or temporal context [6], or both [7]. These deep learning architectures have found significant performance advantages over context-agnostic prediction models. In context-agnostic prediction, observations of $x_{i}$ at location $i$ are used to predict $y_{i}$ at the same location. Instead, modern deep learning models leverage observations made at spatial and or spatiotemporal context neighborhood to enhance estimations.
While advances in deep learning excel at leveraging spatial information, optimal ways of leveraging geographic location information remain underexplored. Note that several methods are proposed to transform geographic (or projected) coordinates into features for deep learning [8], however, regardless of (and beyond) the transformations used, the general higher-level approach in incorporating location into deep learning, which is still an active area of research, is our focus here.
It is crucial that we first differentiate between spatial and geographic. In the context of a Convolutional Neural Network (CNN), for example, the network learns to identify patterns within the image coordinates system, for instance around point $(x,y)$ in image coordinates, without developing an understanding of geographic location $s_{i}$ corresponding to latitude $\varphi$ and longitude $\lambda$ mapped to $(x,y)$. However, how to best incorporate location (i.e., latitude $\varphi$ and longitude $\lambda$) to maximize generalizability remains an open question, with little research quantifying the generalizability impacts of doing so using different methods on temporally-dynamic estimation scenarios. This is partly due to the gap between studies applying deep learning in geospatial sciences and environmental monitoring, where predictive performance is measured on training and testing sets with minimal interrogation of the black box nature of deep learning, and the scarcity of fundamental studies quantifying the behavior of (geo)location information within deep learning frameworks. Investigating the role of location in deep learning-based estimation is even more important given the emergence of ambitious location encoders [9, 10, 11], which attempt to summarize all relevant information at location $s$ into an embedding vector $e_{s}$, suitable for ingestion by deep learning algorithms, for improved predictive performance, and more importantly, geographic generalizability.
Prior research has made significant progress in identifying strategies to incorporate spatial dependence [12] or geographic information into deep learning, including direct use of raw coordinates, hand-engineered spatial features, and learned location embeddings. Mai et al. [8] provide a comprehensive survey of coordinate transformation methods—such as sinusoidal encodings, spatial tile embeddings, and kernel-based representations. Separately, contrastive learning frameworks such as Contrastive Spatial Pre-training (CSP) [9], SatCLIP [10], and GeoCLIP [11] have adapted CLIP-style Contrastive Language-Image Pre-trainig architectures [13] to align image and location representations in a shared latent space (dubbed Contrastive Location-Image Pre-training), enabling pretrained models to distill contextual geographic knowledge. However, evaluations of these location encoders have been limited to static tasks such as species distribution or primitive tasks such as predicting long-term climate averages, where temporal dynamics and real-time variability are less critical, and real-world application is limited. On the other hand, studies such as [14, 15, 16] have proposed robust spatial validation—such as checkerboard partitioning or distance-aware cross-validation—to more accurately assess geographic generalizability. Yet, the interplay between location encoding strategies and spatial evaluation methodology remains underexplored, particularly for dynamic, high-resolution prediction tasks.
In this paper, we take a step towards closing this gap by quantifying and analyzing the impact of incorporating geolocation information of locations $s_{i}$ with latitude $\varphi$ and longitude $\lambda$ in deep learning within a real-world temporally-dynamic application context of estimating daily surface-level PM2.5 from remote sensing imagery. We first offer an in-depth discussion of different approaches to incorporating geolocation into deep learning, including simple featurization to leveraging the state-of-the-art location encoders. We then characterize the domain of surface-level air pollution estimation, including a brief overview and justification for selecting this application domain as an appropriate test-bed for this study. We then present a series of experiments and results from two complementary aspects of predictive (i.e., estimation) performance and geographic generalizability using Within-Region (WR) and Out-of-Region (OoR) evaluation scenarios. We present an ablation study to compare alternatives, and complement the experimental results with qualitative analysis of estimations in the contiguous United States. We characterize the findings and suggest directions for future research.
While the primary contribution of this paper is centered on methods of incorporating geolocation into deep learning, our work also contributes to the literature on surface-level air pollution estimation. Recent advances in surface-level PM2.5 estimation increasingly leverage deep learning architectures that capture spatiotemporal contextual observations, as reviewed in depth by a recent survey [17]. Architectures combining convolutional and recurrent layers, such as CNN-LSTM models [18], have shown improved accuracy by jointly modeling spatial features and temporal dynamics. Graph-based models, including Graph Neural Networks (GNNs) and spatiotemporal Graph Convolutional Networks (GCNs), which represent spatial relationships as graphs, have been used primarily in the context of short-term forecasting at monitoring stations rather than full-coverage spatial surface level estimation [19, 20]. Additionally, attention-based models capture directional flow and long-range dependencies, further improving estimation or forecasting performance [21, 22]. These research efforts show a collective movement and growing emphasis on learning both spatial and temporal context to enhance the robustness and generalizability of PM2.5 estimation frameworks.
Unlike existing PM2.5 estimation established products that rely on context-agnostic regression with raw geographic coordinates [23], treat spatial context as static features like land use or proximity to emission sources [24], or more recent research that leverage spatiotemporal context [17], our current study advances the field by explicitly interrogating the role of geolocation features in deep learning for dynamic, high-resolution air pollution estimation. Our prior work introduced a state-of-the-art Bi-LSTM with attention architecture that demonstrated superior performance, particularly during high-pollution events such as wildfires, by incorporating temporal inputs, including aerosol and meteorological data, and by integrating wildfire smoke density as a predictive covariate [25, 26, 27]. This paper extends that study and evaluates and compares multiple geolocation integration strategies within the modeling framework. By evaluating both within-region (WR) and out-of-region (OoR) performance using several spatial partitioning schemes, we move beyond conventional random or simpler spatiotemporal mix validation schemes common in air pollution studies. This explicit attention to geographic generalizability and the evaluation of location encoder representations in a dynamic prediction task distinguishes our work from prior efforts, including Di et al. [23], Wei et al. [28], and even our own earlier study [25], which did not examine location feature representations or spatial generalization as central research questions.
2 Materials and Methods
2.1 Formalizing Spatiotemporal Context
To formalize spatial and temporal context in this manuscript, consider the following:
2.1.1 Spatial Context
@@
Given a set of spatial locations $S=\{s_{1},s_{2},\dots,s_{n}\}$, where $s_{i}$ represents the coordinates of location $i$, the observation $x_{i}$ at location $s_{i}$ is influenced by its spatial neighborhood $N_{s}(s_{i})=\{s_{j}:d(s_{i},s_{j})\leq r,j\neq i\}$, where $d(\cdot,\cdot)$ can be a distance metric and $r$ is a threshold radius (or another definition of connectivity or neighborhood). A spatial prediction model can be formulated as:
@@
where $f(\cdot)$ is the prediction function learned by the model. Among such models, Convolutional Neural Networks (CNNs) [29] are notable. CNNs operate on data arranged on a regular spatial grid, where features are extracted from each pixel in relation to its local neighborhood defined by a convolution window. This architecture has been extensively used in geospatial and remote sensing applications, spanning tasks in both land [30] and ocean remote sensing [31].
2.1.2 Temporal Context
For temporal sequences, consider a discrete time index $T=\{t_{1},t_{2},\dots,t_{T}\}$. The observation $x_{i}^{t}$ at location $i$ and time $t$ is influenced by its past observations $\{x_{i}^{t-k},\dots,x_{i}^{t-1}\}$ for some lag $k$. A temporal prediction model is represented as:
@@ 2.1.2 Temporal Context
2.1.3 Spatiotemporal Context
Combining both spatial and temporal dimensions, the prediction at location $i$ and time $t$ depends on the spatiotemporal neighborhood:
@@ 2.1.3 Spatiotemporal Context
In this family of models, ConvLSTMs [34] capture spatiotemporal context and have gained adoption in environmental monitoring [35]; Transformers [4] are being employed for remote sensing and geospatial applications [36], and spatiotemporal graph neural networks have shown promise in spatial epidemiology [16], or even in remote sensing for building change detection [37]. The taxonomy above provides a non-exclusive overview of deep learning approaches incorporated in geospatial data science; for instance, sequence-to-sequence methods are not discussed here, but the reader can refer to [38].
@@
The intention is to highlight that the current revolutionary impact of deep learning methods is in large due to their ability to successfully capture spatiotemporal context, and how there is potential for further enhancement if (geo)location information is preserved within the spatiotemporal context.
2.2 Geolocation and Deep Learning
Deep learning is capable of capturing spatial and temporal context; however, methods to incorporate explicit geographic location remains an active area of research. In this section, we examine approaches to integrating geolocation—specifically latitude $\varphi$ and longitude $\lambda$—into deep learning for geospatial estimation tasks. For clarity, we focus on the role of location $s_{i}$ in augmenting a set of observed predictive variables $x_{1}$ through $x_{m}$, with the goal of improving both estimation accuracy and geographic generalizability. The methods discussed here can be extended to temporal and spatiotemporal context models, including the framework used in our experiments.
2.2.1 Approach 1: No Geolocation Feature
@@
2.2.2 Approach 2: Geolocation as Geographic Coordinates Features
@@
While a variety of coordinate transformation strategies have been surveyed before [8], we adopt the sinusoidal encoding as a representative example to evaluate the impact of geolocation features. In our contiguous U.S. study region, where latitude and longitude vary smoothly without discontinuity, the choice of transformation method has limited impact on generalizability.
Including $\varphi$ and $\lambda$ as input features enables deep learning models to learn complex, location-dependent mappings to the target variable $y$. While this can improve performance in training regions, it also risks overfitting to location-specific patterns. For example, the winning team of the AutoICE challenge [39] reportedly [40] used geolocation features to enhance performance in a test region that overlapped with training data (the test images in that competition were situated within the spatial extent of the training data). Put differently, geographic overlap can inflate apparent generalization when location is used as a feature.
@@
2.2.3 Approach 3: Geolocation as Pre-learned Location Encodings
@@
A general formulation of the positional encoder (based on a fourier expansion) is:
@@
2.3 Experiments
2.3.1 Domain Characterization: Estimating Surface-level PM2.5
Estimating surface-level air pollution, and in particular, PM2.5 is important to quantify for public health studies. Prolonged exposure to high concentrations of PM2.5 has been associated with respiratory and cardiovascular diseases, and premature mortality [41]. Accurate, high-resolution spatiotemporal PM2.5 estimates have several important uses in policy-making and public health. Estimates covering recent years are used to quantify PM2.5 human exposure in epidemiological or clinical cohort studies evaluating the impact of air pollution on specific health outcomes, which in the U.S. feed directly into regulatory decisions under the Clean Air Act. Predictions covering future days, on the other hand, are used for public health warnings, while predictions encompassing future years are used to evaluate the effects of regulatory or climatic changes.
In the United States, the Environmental Protection Agency (EPA) has deployed a series of carefully-calibrated Air Quality System (AQS) monitors to measure the spatially and temporally variable concentrations of PM2.5. However, public health studies require estimation of surface-level pollution at the addresses of study cohort members, while the AQS stations may be far away and spatially sparse. On the other hand, satellite observations can be used to estimate particulates suspended in columns of air, and therefore, the predictive task is to use satellite observations and ancillary data to derive surface-level (rather than total column) concentrations of PM2.5. The machine learning target is the surface-level PM2.5 levels measured by AQS sensors. Satellite observations are often leveraged for this purpose. Most notably, Aerosol Optical Depth (AOD), which quantifies the extinction of solar radiation by aerosol particles integrated throughout the atmospheric column, provides valuable large-scale coverage. However, it does not directly measure surface-level concentrations and exhibits a complex, nonlinear relationship with PM2.5 due to meteorological influences, aerosol type, and vertical distribution.
@@ 2.3 Experiments
2.3.2 Model Architecture
Our base model for PM2.5 concentration prediction (Fig.1(a)) follows a Bidirectional Long Short-Term Memory (Bi-LSTM) [6, 42, 43] network with Luong Attention [44], as detailed in [25]. In brief, a 21-day window of time-series multi-variate input is fed into a Bi-LSTM to capture forward and backward dependencies in PM2.5-related features. This 21-day window allows capturing of potential weeks-long persistence of PM2.5 episodes which previous research has identified [45], giving the model ample temporal context while keeping computational demands manageable. Exploratory modeling confirmed that this length achieved the lowest error, with longer windows providing no additional improvement.
@@ 2.3.2 Model Architecture
Inputs to the model include satellite-derived Aerosol Optical Depth (AOD) [46], meteorological variables (e.g., temperature, precipitation, wind direction/speed) [47, 48], wildfire smoke density [27, 26], elevation [49], Normalized Difference Vegetation Index (NDVI) [50], K-Nearest Neighbors Inverse Distance Weighted (IDW) PM2.5 measurements to incorporate ground-based spatial context, as well as temporal encodings (sine and cosine of (Day of the Year, Month of the Year), and Year). Each predictor feature was re-projected to a common 1 km grid (MODIS Sinusoidal) and, when necessary, resampled to match. Daily variables were matched by calendar date, while coarser products—such as the 16-day MODIS NDVI—were assigned to each estimation day using the nearest-available composite. The input datasets (predictors) and target are summarized in Table 1.
Table 1: Datasets, features (predictors), and target used in this study.
Category Variables / Features Source (Product) Spatial Res. Temporal Res. #
Target Surface-level PM2.5 concentration U.S. EPA AQS Station (point) Daily 1
[…]
Meteorology: dayl, prcp, srad, tmax, tmin, vp Daymet 1 km Daily 6
Meteorology: wind direction (th), wind speed (vs) gridMET (1/24) $\sim$4 km Daily 2
Wildfire Smoke Density (WSD) NOAA HMS Smoke Polygon Daily 1
Elevation GMTED2010 1 km Static 1
NDVI (combined) MODIS MCD43A4 500 m 16-day 1
[…]
Temporal encodings: $\sin/\cos$ (DOY, Month), Year Derived N/A Daily 5
Geolocation (variants) Latitude, Longitude Derived (coordinates) N/A Static 2
$\sin/\cos$ of lat, lon Derived (coordinates) N/A Static 4
GeoCLIP location embedding Pretrained location encoder N/A Static $512\text{-}D$
[…]
@@
These features are selected to capture the primary drivers of surface-level PM2.5 concentrations. AOD is satellite-observed columnar aerosol loadings, which, though not a direct measure of surface-level concentrations, is predictive of surface-level pollution under certain atmospheric conditions. Meteorological variables influence pollutant dispersion, chemical transformation, and accumulation. Wildfire smoke density directly relates to episodic spikes in PM2.5 primarily due to biomass combustion. NDVI provides information on vegetative cover, which can modulate emissions and pollutant deposition. Elevation influences local meteorology and pollutant trapping, while the KNN-IDW interpolated PM2.5 input feature reflects observed surface-level spatial trends from monitoring stations. The supervised target in this study is the observed daily PM2.5{2.5} concentration measured at AQS monitoring stations; the KNN–IDW PM2.5 is used only as an auxiliary predictor to convey local spatial context. Note that the IDW-interpolated PM2.5 excludes self-measurements of a monitoring station during training or evaluation, and instead, uses only the 9 nearest neighboring stations’ ground-level measurements to calculate the interpolated values. This is to ensure there is no leakage in training data, and that estimates can be made for any point on the surface. The interpolated feature represents only a distance-weighted neighborhood average value and does not encode the spatial identity or exact location of any monitoring station. Because the Bi-LSTM model has no spatial receptive field and processes each sample independently, it cannot memorize or infer the true values at held-out test stations. Also, note that because inverse-distance weighting rapidly decreases the influence of distant stations, this feature reflects primarily the immediate spatial context rather than a continuous interpolated surface. The KNN–IDW feature may be less informative in areas with sparse monitoring; however, the Bi-LSTM model uses it jointly with other predictors—such as AOD, meteorology, and vegetation—that more directly capture the pollution level in those regions.
Because AOD is a key predictor of surface PM2.5 and a time-series of it is fed to the model for every pixel, missingness caused by cloud or snow cover were first imputed with a lightweight random forest regressor trained on the same set of predictors (excluding the KNN-IDW PM2.5 and the target itself), as detailed in [51]. After this step, a single masking layer flags any residual missing entries across all predictors, enabling the Bi-LSTM to downweight them during training rather than treat them as observed values [52].
@@
Layer normalization and dropout [53] are applied after each Bi-LSTM layer to stabilize training and mitigate overfitting. The Luong attention mechanism [44] then learns to focus on the most informative timesteps within the 21-day window by computing a context vector as a weighted average of the hidden states. The representation is then passed through a fully connected layer to make the final prediction for PM2.5. We refer the reader to [25] for detailed information on data processing, implementation, long-term evaluation of the model and comparison to multiple other baselines and benchmarks. What is different in this manuscript is the ways in which geolocation features are fed to the model, and the way performance is evaluated under regional conditions. Specifically, we use several configurations in regional partitioning for creating training and testing splits as detailed in the Section 3, and for each configuration, train the model with four general variations—differing in the way geolocation is incorporated:
@@
Figure 1: Model architecture for estimating surface-level PM2.5 concentrations enhanced with a) no geolocation info, or geolocation appended as geographic coordinates features, and b) geolocation embeddings vector extracted from the pretrained location encoder and fused through the Hadamard product operation.
@@
    No geolocation input: The model does not receive geolocation $(\varphi,\lambda)$ information, similar to Approach 1 described in section 2.2, forcing the model to purely rely on time-series observations for estimating the target.
@@
    Direct geolocation input: We append the raw latitude and longitude values $(\varphi,\lambda)$ to the time-series feature vector at each timestep, similar to Approach 2 described in section 2.2, allowing the model to rely on both the time series observations as well as geolocation for estimating the target.
@@
    Sinusoidal transformation of geographic coordinates: Also similar to Approach 2 in section 2.2, but with a transformation applied first: instead of passing latitude and longitude values directly in angular units, we apply sinusoidal transformations to preserve their cyclical properties:
@@
    Fusing location encoder embeddings: We integrate pre-learned location embeddings from the pretrained location encoder GeoCLIP [11]. GeoCLIP is a location encoder trained using a CLIP-style contrastive learning framework to align geotagged Flickr images with geographic coordinates (section 2.2.3 above covers details of this approach). Its image encoder is a Vision Transformer (ViT-B/16), while the positional encoder is a Fourier feature mapping of latitude and longitude, capturing multi-scale spatial patterns. During pretraining by the original developers [11], GeoCLIP was trained on over 1.2 million image-location pairs globally sampled from the YFCC100M dataset, with relatively denser coverage in urbanized and photo-rich regions, which is important for PM2.5 estimation in areas more prone to generate emissions. We will contextualize GeoCLIP against other available location encoders in section 3.3.